Niobium with an atomic weight of 92.90 u is characterized by a bcc lattice structure with a lattice constant of 3.30 Å. Calculate the density.


1
Expert's answer
2021-02-17T11:09:26-0500

The density of niobium can be found as follows:


"\rho=\dfrac{nA}{a^3N_A},""\rho=\dfrac{2\cdot92.9\ amu\cdot\dfrac{g}{mol}}{(3.3\cdot10^{-8}\ cm)^3\cdot6.023\cdot10^{23}\ mol^{-1}}=8.58\ \dfrac{g}{cm^3}."